OpenStudy (anonymous):

please help to solve and i will fan and give a medal Y" + x y' + y= 0 when y(0)=3 and y'(0)=1

OpenStudy (rational):

what have you tried so far ?

OpenStudy (anonymous):

i have tried to solve it but i have not found a good book for it

OpenStudy (rational):

you may find similar example problems at below site useful http://tutorial.math.lamar.edu/Classes/DE/SeriesSolutions.aspx

OpenStudy (anonymous):

its too complex for me and its an exam question

OpenStudy (anonymous):

please help me

OpenStudy (rational):

suppose the solution can be represented as power series : \[y = \sum\limits_{n=0}^{\infty} a_nx^n\]

OpenStudy (anonymous):

please continue sir

OpenStudy (rational):

find \(y'\) and \(y''\) then plug them into the given differential equation

OpenStudy (anonymous):

ok but can you please help with solution.. its easier if it is solved and i follow thesame step please

OpenStudy (rational):

\(y = \sum\limits_{n=0}^{\infty} a_nx^n\) \(y' = \sum\limits_{n=1}^{\infty} na_nx^{n-1}\) \(y'' = \sum\limits_{n=2}^{\infty} n(n-1)a_nx^{n-2}\) the given differential equation : \[y" + x y' + y= 0\] becomes : \[ \sum\limits_{n=2}^{\infty} n(n-1)a_nx^{n-2} + x \sum\limits_{n=1}^{\infty} na_nx^{n-1}+ \sum\limits_{n=0}^{\infty} a_nx^n= 0\]

OpenStudy (anonymous):

ok that is the y" ,y' and y in the equation above. thanks sir please continue

OpenStudy (rational):

Next the trick is to make the exponent of "x" equal in all terms : \[\sum\limits_{n=2}^{\infty} n(n-1)a_nx^{n-2} + \color{red}{x} \sum\limits_{n=1}^{\infty} na_n\color{Red}{x^{n-1}}+ \sum\limits_{n=0}^{\infty} a_nx^n= 0\] Multiplying them gives \[\sum\limits_{n=2}^{\infty} n(n-1)a_nx^{n-2} + \sum\limits_{n=1}^{\infty} na_n\color{Red}{x^{n}}+ \sum\limits_{n=0}^{\infty} a_nx^n= 0\]

OpenStudy (rational):

Notice the exponent in first term is \(x^{n-2}\), lets see how to make it \(n\) ?

OpenStudy (anonymous):

ok prof

OpenStudy (anonymous):

please ride on

OpenStudy (rational):

substitute \(n-2 = k\) then \(n = k+2\) \[\sum\limits_{k=0}^{\infty} (k+2)(k+1)a_{k+2}x^{k} + \sum\limits_{n=1}^{\infty} na_n\color{Red}{x^{n}}+ \sum\limits_{n=0}^{\infty} a_nx^n= 0\] which is same as \[\sum\limits_{n=0}^{\infty} (n+2)(n+1)a_{n+2}x^{n} + \sum\limits_{n=1}^{\infty} na_n\color{Red}{x^{n}}+ \sum\limits_{n=0}^{\infty} a_nx^n= 0\] yes ?

OpenStudy (anonymous):

sir please note that the intial conditions given are y(0)=1 and y'(0)=0 please forgive me for my mistake . continue please

OpenStudy (rational):

we will use initial conditions shortly, did u get how we managed to get all the terms to have same exponent ?

OpenStudy (anonymous):

yes very well sir

OpenStudy (rational):

\[\sum\limits_{n=0}^{\infty} (n+2)(n+1)a_{n+2}x^{n} + \sum\limits_{n=1}^{\infty} na_n\color{Red}{x^{n}}+ \sum\limits_{n=0}^{\infty} a_nx^n= 0\] which is same as \[\sum\limits_{n=0}^{\infty} \left[ (n+2)(n+1)a_{n+2}x^{n} + na_n\color{Red}{x^{n}}+ a_nx^n\right]= 0\]

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

what next sir

OpenStudy (anonymous):

sir please continue and complete the solution . my system battery is dying . i will be viwing it with my phone but might not type any thing .... please hope you will complete it for me. GOD bless you....this is the best site

OpenStudy (rational):

\[\sum\limits_{n=0}^{\infty} \left[ (n+2)(n+1)a_{n+2}x^{n} + na_n\color{Red}{x^{n}}+ a_nx^n\right]= 0\] \[\sum\limits_{n=0}^{\infty} \left[ (n+2)(n+1)a_{n+2} + (n+1)a_n\right] x^n = 0\] set the coefficient to 0 and solve the recurrence relation \[ (n+2)(n+1)a_{n+2} + (n+1)a_n = 0\]

OpenStudy (anonymous):

ok. i get this step sir. what next

OpenStudy (anonymous):

please complete it for me sir. i will look at how you solved every thing and learn. GOD bless

OpenStudy (rational):

isolating \(a_{n+2}\) gives : \[a_{n+2} = -\dfrac{(n+1)a_n}{(n+2)(n+1)}\] use the initial conditions and see if you can sovle above

OpenStudy (anonymous):

Non-power series approach: Notice that you have two derivatives in the ODE: \[\frac{d}{dx}[y']=y''\quad\text{and}\quad\frac{d}{dx}[xy]=xy'+y\] So \[\begin{align*} y''+xy'+y&=0\\\\ \frac{d}{dx}[y'+xy]&=0\\\\ y'+xy&=C_1 \end{align*}\] which is linear in \(y\).

OpenStudy (anonymous):

@ rational, i understood all you did here but i could not finish it. please continue and finish it for me. i have less time before my exam. please help me on this . thanks and God bless

OpenStudy (anonymous):

please do not forget i made a modification on the initial conditions given